OpenStudy (anonymous):

d=1.5 square root of h. solve for h

OpenStudy (oaktree):

First divide by 1.5, and then square both sides.

OpenStudy (anonymous):

\[d=1.5 \sqrt{h}\]

OpenStudy (oaktree):

Right...now divide by 1.5

OpenStudy (oaktree):

\[d/1.5=\sqrt(h)\]

OpenStudy (oaktree):

Then square both sides to get rid of the square root.

OpenStudy (oaktree):

\[d^2/2.25=h\]

OpenStudy (anonymous):

thank you